HelpMePlz:

A confectioner is going to mix candy worth $3.90 per pound with candy worth $2.50 per pound to obtain 70 pounds of candy worth $3.30 per pound. How many pounds of each kind should she use?

Champion:

say you use 20 pounds of the $3.90 candy. how many pounds of the $2.50 would you use?

HelpMePlz:

idk

Champion:

hint, the total is 70 pounds, you used 20

HelpMePlz:

oh ok 50

Champion:

right , and what would the cost of the $3.90 candy be total?

HelpMePlz:

um 3.90 x 20 ?

Champion:

yeah right aka $78

HelpMePlz:

oh ok and the 50 pounds at $2.50 would cost $125 right?

Champion:

yeah, 2.50×50=125 so what would the total be?

HelpMePlz:

add right? 78+125=203

Champion:

yes, and 203 at 70 pounds would be what ?

HelpMePlz:

203/70=2.9?

HelpMePlz:

so what do we do now?

Champion:

try again, this time with variables say you use x amount of $3.90 candy, how much of the $3.90 candy do you use?

HelpMePlz:

??

Champion:

how did you get the 50?

HelpMePlz:

oh subtract 70−x right?

Champion:

yes subtract now the x pounds of $3.90 candy cost 3.90×x how much does the other candy cost?

HelpMePlz:

must be 2.50×(70−x) ?

Champion:

right, and the total has to be 70×3.30 so set 3.9x+2.5×(70−x)=70×3.3 and solve for x

HelpMePlz:

ooh ok

HelpMePlz:

I get it, thanks!